Enough of a monoprotic weak acid is dissolved in water to produce a 0.0113 M solution. The pH of the resulting solution is 2.64 . Calculate the Ka for the acid.
Answer #1


use:
pH = -log [H+]
2.64 = -log [H+]
[H+] = 2.291*10^-3 M
HA dissociates as:

HA          ----->     H+   + A-
1.13*10^-2                 0         0
1.13*10^-2-x               x         x


Ka = [H+][A-]/[HA]
Ka = x*x/(c-x)
Ka = 2.291*10^-3*2.291*10^-3/(0.0113-2.291*10^-3)
Ka = 5.825*10^-4
Answer: 5.8*10^-4